# sample.bib

### v2

@@ Abstract
Because of the rudimentary reporting methods and general lack of documentation, the creation of a severe weather database within the Philippines has been difficult yet relevant target for climatology purposes and historical interest. Previous online severe weather documentation i.e. of tornadoes, waterspouts, and hail events, has also often been few, inconsistent, inactive, or is now completely decommissioned. Several countries or continents support severe weather information through either government-sponsored or independent organizations. For this work, Project SWAP stands as a collaborative exercise, with clear data attribution and open avenues for augmentation, and the creation of a common data model to store the phenomenon’s information will assist in maintaining and updating the aforementioned online archive in the Philippines. This paper presents the methods necessary for creating the SWAP database, provide broader climatological analysis of spatio–temporal patterns in severe weather occurrence within the Philippine context, and outline potential use cases for the data. We also highlight the project’s current limitations as is to any other existing and far larger database and emphasize the need for understanding these events’ and their mesoscale environments, inline to the current severe weather climatologies across the globe.
@@ 2.1 Data Sources: Tornado, Hail, and Waterspout database
Information not included in any major available dataset was obtained from other reliable sources as indicated in De Coning and Adam [84], which were used both to add new tornadoes and fill in missing data for existing ones. Doing so is quite time-consuming, as a vast number of different sources are involved, and the effort to find and parse them is continuing. The non-official sources include eyewitness reports, media news/newspapers [^4] Most of these newspapers are available at the National Library of the Philippines (NLP) upon request. Same goes to other official sources such as those from DOST-PAGASA whose documents are very difficult to find in the internet., and information from social media platforms (e.g. Twitter, Facebook, and YouTube). Such documentary data were exhaustively evaluated to identify fake events. The updated database includes, as far as possible, information about the type of event, date, hour, location (latitude-longitude coordinates), tags e.g. tornadoes, hails, elevated, high-based etc., photographs and videos through the documentation column, and other relevant information. To regulate the process, every tornado in the database was given a mandatory “Source” attribute which either referred to one of our major datasets or another source (usually a web link). We designate these sources whether they are still functioning (Active), preserved somewhere else like a library (Preserved), or no longer accessible or maintained (Inactive).
@@ 3.1 Limitation of Project SWAP
    Number of SWEs. Many SWEs go unreported or unrecorded, especially in areas or time periods with less infrastructure or lower population density [34, 35]. This is reflected, for example, prior to 2010s including the years 2010 and 2016, were only 2 and 8 cases were tallied to the archive, respectively. We are not sure why it is difficult to look for SWEs way back on Year 2010, but the lack of cases in Year 2016 was likely due to the August 14 Manila Tornado that occurred on the Metro, making it a standout signature, while the rest of the SWEs were concealed by the alias.
@@ 3.2 Temporal Distribution of Severe Weather Activity
Figure 2: (a) Monthly and (b) seasonal distribution of documented tornadoes, hail, waterspouts, and total severe weather activity in Philippines.
@@
In summary, the severe weather season kicks off in spring (March, April, and May, MAM), reaches its peak in summer (JJA) along with reported hail decreasing, accompanied by the winding down of SWEs in autumn (SON), and finishes in winter (DJF) (Fig. 2b). There is an evident difference between tornadoes, commonly reported in spring and summer, and waterspouts, which have their active period between summer and autumn. It is inferred that waterspouts are notably influenced by tropical cyclone activity, principally along the eastern coast of the Pacific Ocean. On the other hand, tornadoes are more related to instability conditions derived from the pass of easterly waves, moisture advection, favorable wind profile, and the shear-line activity [^6] Also known as Tail-end of a Frontal System, usually a Cold Front..
@@ 3.3 Spatial Distribution of Severe Weather Activity
Figure 4a shows the geographic distribution of all documented SWEs across the archipelago. The highest tornadic and hail activity is registered in Greater Metro Manila (GMM) area encompassing Southern Luzon, National Capital Region (NCR), and Region III (Central Luzon). While scattered area of waterspout events were seen in along the coast of Southern Luzon and even at Laguna Lake, notably with the quadruple spouts back in May 2020 at height of pandemic. On the other hand, a large swath of waterspout events coinciding with few tornadic events was recorded in the Regions VI and VII (Western and Central Visayas). Lastly, another area where significant number of tornadoes were recorded was evident along BARMM (Bangsamoro Autonomous Region in Muslim Mindanao) and Region XII (SOCKSARGEN) with North and South Cotabato, Sultan Kudarat, and Maguindanao as its favored corridor for severe weather activity, making it 2nd on the highest tornadic activity. Notably, a regional tornado outbreak occurred in South Cotabato at the afternoon of September 30, 2009 impacting several municipalities, including Koronadal City.
@@
Figures 6a, 6b, and 6c displays a kernel density estimation made from sci-kit learn Python package on 11 km x 11 km grid. A clear and well-defined hotspot for all hazards threat (tornadoes, hails, and waterspouts) with bimodal distribution was latched in the GMM. The greatest likelihood for severe weather at the upper-portion of Metro Manila-Pampanga area and another at Pangasinan, while another hail hotspot was denoted along Benguet region (Fig. 6a,b). A total of 137 tornado and 105 hail reports have been documented over this area i.e. $\geq$40and $\geq$70%, respectively of the total within SWAP DR2. Some of the notable events include the first recorded tornado in the Philippines back on June 1968 by Grazulis [83], a potential EF2 tornado that tore parts of Bulakan, Bulacan on 1998 August, the well-documented EF1 Manila Tornado on 2016 August, the quadruplet waterspouts in Laguna Lake on 2020 May, an significant 8 cm hail in Norzagaray, Bulacan on 2021 August, and yet another Pampanga Tornadoes that impacted towns of Magalang and Arayat back on 2023 June and May of this year, respectively.
@@
Additionally, the region has a historical record of significant tornado events, including three potential EF2 tornadoes. Notably, a killer tornado struck Zamboanga del Sur in 1990 June, resulting in the deaths of up to 30 people in Manukan, despite being outside the main hotspot zone. Other destructive tornadoes were recorded in Lantapan, Bukidnon in 1991 July, a regional tornado outbreak in South Cotabato back on 2009 September, and in Pikit, North Cotabato in 2015 August. The severity and impact of these tornadoes further underscore the vulnerability of the region to SWEs, especially within the identified hotspot zone. However further studies could explore the precise meteorological mechanisms at play, enhancing the understanding of tornadogenesis in this complex terrain.111111as also noted the same for the severe weather hotspot in Luzon.
@@ 4 Conclusion and Recommendation
The monthly distribution shows that the beginning of the most active phase of the severe weather season is in April. This first activity fits with the most active tornado phase in the USA [70]. In May, SWEs have been previously documented across the country, in particular at lower Region I and Region III encompassing GMM area and Southern Luzon, with early termination between September-October. This may be indicating that this portion of the country may be related to similar atmospheric processes that lead to supercell formation and tornadogenesis, especially during Southwest Monsoon periods and the start of TC season in the western Pacific. However, easterly wind setups during months of April and May may also provide the same hazardous meteorological condition. Several studies documented the characteristics and role of tropical activity in rainfall in Philippines [52, 65, 74]. However, the impact of TCs on tornadogenesis and severe weather initiation remains understudied in the Philippine context.
Meanwhile, down south towards the Western and Central Visayas, waterspouts, along some tornadoes in the Panay Islands, are the main show for severe weather activity. The start of increased waterspout and tornado activity is in May followed by scattered and clumps of activity across the island groups of Visayas until it slowly winds down by November. On the other hand, Mindanao, particularly BARRM and Region XII, had a head start in terms of severe weather; mostly tornadic events, from April to November with severe weather season capping off by December across the country. The environments in these locations were generally comprised in accordance to the Philippine Climate Types developed by Coronas [82] and Kintanar [81], but the convective and kinematic setup of these remains to be studied.
@@ Acknowledgments
Project SWAP and its SWAP DR2 is available and distributed on Zenodo. Sounding profiles will be available soon on the aforementioned Zenodo page. The Digital Elevation Model (DEM) is from SRTM15Plus distributed and available on OpenTopography. The population density is available through World Population Hub. Finally, the lightning data from the TRMM is accessible through NASA Earth Data. All of these data is distributed under Creative Commons CC-BY 4.0. Proper attribution is required for these datasets.

### v3

@@ Abstract
Because of the rudimentary reporting methods and general lack of documentation, the creation of a severe weather database within the Philippines has been difficult yet relevant target for climatology purposes and historical interest. Previous online severe weather documentation i.e. of tornadoes, waterspouts, and hail events, has also often been few, inconsistent, inactive, or is now completely decommissioned. Several countries or continents support severe weather information through either government-sponsored or independent organizations. For this work, Project SWAP stands as a collaborative exercise, with clear data attribution and open avenues for augmentation, and the creation of a common data model to store the phenomenon’s information will assist in maintaining and updating the aforementioned online archive in the Philippines. This paper presents the methods necessary for creating the SWAP database, provide broader climatological analysis of spatio–temporal patterns in severe weather occurrence within the Philippine context, and outline potential use cases for the data. We also highlight the project’s current limitations as is to any other existing and far larger database, and emphasize the need for understanding these events’ and their mesoscale environments, inline to the current severe weather climatologies across the globe.
@@ 2.1 Data Sources: Tornado, Hail, and Waterspout database
Information not included in any major available dataset was obtained from other reliable sources as indicated in De Coning and Adam [84], which were used both to add new tornadoes and fill in missing data for existing ones. Doing so is quite time-consuming, as a vast number of different sources are involved, and the effort to find and parse them is continuing. The non-official sources include eyewitness reports, media news/newspapers [^4] Most of these newspapers are available at the National Library of the Philippines (NLP) upon request. Same goes to other official sources such as those from DOST-PAGASA whose documents are very difficult to find in the internet., and information from social media platforms (e.g. Twitter, Facebook, and YouTube). Such documentary data were exhaustively evaluated to identify fake events. The updated database includes, as far as possible, information about the type of event, date, hour, location (latitude-longitude coordinates), tags e.g. tornadoes, hails, elevated, high-based etc., photographs and videos through the documentation column, and other relevant information. To regulate the process, every SWEs in the database was given a mandatory “Source” attribute which either referred to one of our major datasets or another source (usually a web link). We designate these sources whether they are still functioning (Active), preserved somewhere else like a library (Preserved), or no longer accessible or maintained (Inactive).
@@ 3.1 Limitation of Project SWAP
    Number of SWEs. Many SWEs go unreported or unrecorded, especially in areas or time periods with less infrastructure or lower population density [34, 35]. This is reflected, for example, prior to 2010s including the years 2010 and 2016, were only 2 and 8 cases were tallied to the archive, respectively. We are not sure why it is difficult to look for SWEs way back on Year 2010, but the lack of cases in Year 2016 was likely due to the 14 August 2016 Manila Tornado that occurred on the Metro, making it a standout signature, while the rest of the SWEs were concealed by the alias.
@@ 3.2 Temporal Distribution of Severe Weather Activity
Figure 2: (a) Monthly and (b) seasonal distribution of documented tornadoes, hail, waterspouts, and total severe weather activity in Philippines. Averages along the severe weather season is denoted by the solid colored lines.
@@
In summary, the severe weather season kicks off in spring (March, April, and May, MAM), reaches its peak in summer (JJA) along with reported hail decreasing, accompanied by the gradual winding down of SWEs in autumn (SON), and finishes in winter (DJF) (Fig. 2b). On average, 225 SWEs were tallied over the span of the severe weather season, with 120 tornadoes, 55 hail events, and 50 waterspouts displayed along with the colored lines. There is an evident difference between tornadoes, commonly reported in spring and summer, and waterspouts, which have their active period between summer and autumn. It is inferred that waterspouts are notably influenced by tropical cyclone activity, principally along the eastern coast of the Pacific Ocean. On the other hand, tornadoes are more related to instability conditions derived from the pass of easterly waves, moisture advection, favorable wind profile, and the shear-line activity [^6] Also known as Tail-end of a Frontal System, usually a Cold Front..
@@ 3.3 Spatial Distribution of Severe Weather Activity
Figure 4a shows the geographic distribution of all documented SWEs across the archipelago. The highest tornadic and hail activity is registered in Greater Metro Manila (GMM) area encompassing Southern Luzon, National Capital Region (NCR), and Region III (Central Luzon). While scattered area of waterspout events were seen in along the coast of Southern Luzon and even at Laguna Lake, notably with the quadruple spouts back in 2020 May at height of pandemic. On the other hand, a large swath of waterspout events coinciding with few tornadic events was recorded in the Regions VI and VII (Western and Central Visayas). Lastly, another area where significant number of tornadoes were recorded was evident along BARMM (Bangsamoro Autonomous Region in Muslim Mindanao) and Region XII (SOCKSARGEN) with North and South Cotabato, Sultan Kudarat, and Maguindanao as its favored corridor for severe weather activity, making it 2nd on the highest tornadic activity. Notably, a regional tornado outbreak occurred in South Cotabato at the afternoon of September 30, 2009 impacting several municipalities, including Koronadal City.
@@
Figures 6a, 6b, and 6c displays a kernel density estimation made from sci-kit learn Python package on 11 km x 11 km grid. A clear and well-defined hotspot for all hazards threat (tornadoes, hails, and waterspouts) with bimodal distribution was latched in the GMM. The greatest likelihood for severe weather at the upper-portion of Metro Manila-Pampanga area and another at Pangasinan, while another hail hotspot was denoted along Benguet region (Fig. 6a and 6b). A total of 137 tornado and 105 hail reports have been documented over this area i.e. $\geq$40and $\geq$70%, respectively of the total within SWAP DR2. Some of the notable events include the first recorded tornado in the Philippines back on 1968 June by Grazulis [83], a potential EF2 tornado that tore parts of Bulakan, Bulacan on 1998 August, the well-documented EF1 Manila Tornado on 2016 August, the quadruplet waterspouts in Laguna Lake on 2020 May, a significant 8 cm hail in Norzagaray, Bulacan on 2021 August, and yet another Pampanga Tornadoes that impacted towns of Magalang and Arayat back on 2023 June and May of this year, respectively.
@@
Additionally, the region has a historical record of significant tornado events, including three potential EF2 tornadoes. Notably, a killer tornado struck Zamboanga del Sur in 1990 June, resulting in the deaths of up to 30 people in Manukan, despite being outside the main hotspot zone. Other destructive tornadoes were recorded in Lantapan, Bukidnon in 1991 July, a regional tornado outbreak in South Cotabato back on 2009 September, and in Pikit, North Cotabato in 2015 August. The severity and impact of these tornadoes further underscore the vulnerability of the region to SWEs, especially within the identified hotspot zone. However, further studies could explore the precise meteorological mechanisms at play by looking at the role of this complex terrain in tornadogenesis.111111as also noted the same for the severe weather hotspot in Luzon.
@@ 4 Conclusion and Recommendation
The monthly distribution shows that the beginning of the most active phase of the severe weather season is in April. This first activity fits with the most active tornado phase in the USA [70]. In May, SWEs have been previously documented across the country, in particular at lower Region I and Region III encompassing GMM area and Southern Luzon, with early and gradual termination between September-October. This may be indicating that this portion of the country may be related to similar atmospheric processes that lead to supercell formation and tornadogenesis, especially during Southwest Monsoon periods and the start of TC season in the western Pacific. However, easterly wind setups during months of April and May may also provide the same hazardous meteorological condition. Several studies documented the characteristics and role of tropical activity in rainfall in Philippines [52, 65, 74]. Although, the impact of TCs on tornadogenesis and severe weather initiation remains understudied in the Philippine context.
Meanwhile, down south towards the Western and Central Visayas, waterspouts, along some tornadoes in the Panay Islands, are the main show for severe weather activity. The start of increased waterspout and tornado activity is in May followed by scattered and clumps of activity across the island groups of Visayas until it slowly winds down by November. On the other hand, Mindanao, particularly BARRM and Region XII, had a head start in terms of severe weather; mostly tornadic events, stretching from April to November with severe weather season capping off by December across the country. The environments in these locations were generally comprised in accordance to the Philippine Climate Types developed by Coronas [82] and Kintanar [81], but the convective and kinematic setup of these remains to be studied.
@@ Acknowledgments
Project SWAP and its SWAP DR2 is available and distributed on Zenodo. The Digital Elevation Model (DEM) is from SRTM15Plus distributed and available on OpenTopography. The population density is available through World Population Hub. Finally, the lightning data from the TRMM is accessible through NASA Earth Data. All of these data is distributed under Creative Commons CC-BY 4.0. Proper attribution is required for these datasets.